OpenStudy (ammarah):

A flag pole is 12 metres tall. An observer looks up at an angle of 25 degrees to see the top of the pole. How far is the observer from the foot of the flag pole?

OpenStudy (ammarah):

@triciaal @jim_thompson5910

OpenStudy (ammarah):



OpenStudy (ammarah):

Is this correct setup? So I did tan 25 degree= 12/x

OpenStudy (triciaal):

i agree

OpenStudy (ammarah):

x-5.59? do i take the inverse?

OpenStudy (triciaal):

no

OpenStudy (triciaal):

you take the inverse when you already have the value and you want the angle

OpenStudy (ammarah):

ohh k so is my answer correct? im still not used to of the calculator with the tangent stuff so please check

OpenStudy (triciaal):

to solve for x 12/x = tan 25 x = 12/tan 25 = 12/ 0.466 = 25.75

OpenStudy (triciaal):